OpenStudy (anonymous):

solve the equation 3|2-5x|=3x+5

OpenStudy (anonymous):

When 2-5x is >0 we have 6-15x=3x+5 which gives x=1/18

OpenStudy (anonymous):

when 2-5x<0 we have -6+15x=3x+5 x=11/12

OpenStudy (anonymous):

does that help

OpenStudy (anonymous):

yes thank you!